solve the function r=241m^-1/4 for m
solve the function w=9.1m^3/4 for m
explain how to use the above answer to express heart rate r as a function of the brain weight w. then find the function

Answer by stanbon(75887)

solve the function r=241m^-1/4 for m
m^(-1/4) = r/241
m = (r/241)^-4)
-------------------------
solve the function w=9.1m^3/4 for m
m^(3/4) = w/9.1
m = [w/9.1]^(4/3)
-------------------------
explain how to use the above answer to express heart rate r as a function of the brain weight w. then find the function
=============================
Substitute for "m" in the "r" equation:
r = 241^(4/3)^(-1/4)
---
The function: r = 241^(-1/3)
Can also be written as follows:
r = 241^(1/3)
====================================
Cheers,
Stan H.